Solving Systems of Equations by Graphing:
I am having alot of trouble solving this equation, I tried to look at the examples to see how it is done, but i don't understand, this is how the example was written and solved :
Solve the system of equations by graphing:
2x + y = 5
x - y = 1
slope intercept form:
2x + y = 5 ---> y = -2x + 5
x - y = 1 ---> y = x - 1
{and then there is a graph drawn which intersects at (2,1)
my question is: How can you draw a graph using y = -2x + 5 and y = x - 1 ?

( i 'd understand if it was y = 2 or y = 3 but this one has variable 'x' in it too)
please help!

Answer by pwac(253)

y=-2x+5 and y=x-1 will give you two separate straight lines on a graph.
To get y= to a number you substitute x with a number and do equation to get a value for y.
Example for your 1st equation y=-2x+5 let x=2 so you get
y=-2(2)+5 which comes out to:y=1
So the point you would plot on graph (2,1)
when x=2,y=1.
so select a range of values for x say from -10 to 10 in steps of two
plug into your equation to get y values and plot on graph.
repeat for other equation.
Unfortunately I do not know how to use graph diagram on this site yet so cannot not show you.
hope this helps.
Pete